two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively . in how many hours will they meet?
Answer by venugopalramana(3286)

two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively . in how many hours will they meet?
RELATIVE VELOCITY WHILE TRAVELING IN OPPOSITE DIRECTIONS =V1+V2=3+4=7 MPH
DISTANCE OF SEPARATION=14
TIME NEEDED TO MEET =14/7 = 2 HRS